Candence has $5,400 to invest. She can invest in bonds that pay 7.5% interest and CDs that pay 6% interest. She wishes to invest some both ways. How much does she invest in each if she earns $384 interest in one year?
Answer by LinnW(1048)

Let x = the amount invested at 6%
x*6% + (5400-x)*7.5% = 384
0.06x + 0.075(5400-x) = 384
0.06x + 405 - 0.07x = 384
-0.01x + 405 = 384
add -405 to each side
-0.01x = -21
divide each side by -0.01
x = 1400
So 1400 at 6% = 84
(5400-1400) or 4000 at 7.5% = 300
Our total interest is 384
So we invest 1400 at 6% and 4000 at 7.5%